a rectangle is twice as long as it is wide. what is the ratio of the wide of the rectangle to its perimeter?
Answer by dfrazzetto(283)


Length = L
Width = W
Perimeter = 2L + 2W
L = 2W
P = 2L + 2W = 2(2W) + 2W = 4W + 2W = 6W
P = 6W; Perimeter expressed in terms of Width

Ratio of width to perimeter:
ratio = W / P = W / 6W
Ratio of width to perimeter = 1/6